Solve (x²D² + 4xD + 2)y= (logx/x)^2,x> 0, where D =d/dx.


1
Expert's answer
2021-08-16T02:19:50-0400

The auxiliary equation is

"m^2+4m+2=0\\
m_{1,\:2}=\frac{-4\pm \sqrt{4^2-4\cdot \:1\cdot \:2}}{2\cdot \:1}\\
m=-2+\sqrt{2},\:m=-2-\sqrt{2}"

Thus the complementary solution is

"y_c= (c_1+xc_2)e^{-x}\\
Let \space y_1= e^{-x} ; y_2= xe^{-x}\\
w= \begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}= \begin{vmatrix}
 e^{-x} & xe^{-x}\\
 -e^{-x} & -xe^{-x}+e^{-x}
\end{vmatrix}= e^{-2x}\\"

By the method of variation of parameter particular solution is

"y_p=-y_1 \int \frac{y_2 f(x)}{w}dx+y_2 \int \frac{y_1 f(x)}{w}dx\\
y_p=x^2e^{-x}[\frac{log x }{2}+ \frac{1}{4}+log x -1]\\
y_p= \frac{x^2e^{-x}}{4}[2log x -3]\\"


Hence the general solution is

"y(x)= (c_1+xc_2)e^{-x}+\frac{x^2e^{-x}}{4}[2log x -3]"